# RGNMR: A Gauss-Newton method for robust matrix completion with theoretical guarantees

### v1

@@ Abstract
Recovering a low rank matrix from a subset of its entries, some of which may be corrupted, is known as the robust matrix completion (RMC) problem. Existing RMC methods have several limitations: they require a relatively large number of observed entries; they may fail under overparametrization, when their assumed rank is higher than the correct one; and many of them fail to recover even mildly ill-conditioned matrices. In this paper we propose a novel RMC method, denoted RGNMR, which overcomes these limitations. RGNMR is a simple factorization-based iterative algorithm, which combines a Gauss–Newton linearization with removal of entries suspected to be outliers. On the theoretical front, we prove that under suitable assumptions, RGNMR is guaranteed exact recovery of the underlying low rank matrix. Our theoretical results improve upon the best currently known for factorization-based methods. On the empirical front, we show via several simulations the advantages of RGNMR over existing RMC methods, and in particular its ability to handle a small number of observed entries, overparameterization of the rank and ill-conditioned matrices.
@@ 1 Introduction
Figure 1: Performance of RMC methods as a function of the number of observed entries measured by the oversampling ratio $\frac{|\Omega|}{r\cdot(n_{1}+n_{2}-r)}$ where $r\cdot(n_{1}+n_{2}-r)$ is the number of degrees of freedom of a rank $r$ matrix. (left) Median rel-RMSE; (right) Failure probability (±$\pm 1.96$ SE). The underlying matrix $L^{*}$ has a condition number $\kappa=2$.
Figure 2: Performance of RMC methods under overparameterization with input rank of $5+i$ for $i\in[0,5]$. (left) Median rel-RMSE; (right) Failure probability (±$\pm 1.96$ SE). The matrix $L^{*}$ has a condition number $\kappa=2$ and the oversampling ratio is $\frac{|\Omega|}{r\cdot(n_{1}+n_{2}-r)}=12$.
Figure 3: Performance of RMC methods as a function of the condition number. (left) Median rel-RMSE; (right) Failure probability (±$\pm 1.96$ SE). The oversampling ratio is $\frac{|\Omega|}{r\cdot(n_{1}+n_{2}-r)}=12$.
Low-rank matrices play a fundamental role in multiple scientific disciplines. As reviewed in Davenport and Romberg, (2016), in various applications, there is a need to recover a low rank matrix from only a subset of its entries. Examples include recommendation systems (Bennett et al.,, 2007), various problems in computer vision (Kennedy et al.,, 2016; Tulyakov et al.,, 2016; Miao and Kou,, 2021), and sensor network localization (Drineas et al.,, 2006). A key challenge in these and other applications is that some of the observed entries may be arbitrarily corrupted outliers.
@@
Over the past decade, the RMC problem was studied from both mathematical, computational and statistical perspectives. Several RMC algorithms have been proposed. In general, to solve the RMC problem the following optimization issues need to be addressed: (i) either promote a low rank solution or strictly enforce it; and (ii) suppress the potentially detrimental effect of the outliers, whose locations are a-priori unknown.
Regarding the first issue, one approach is to incorporate into the optimization objective a low-rank promoting penalty such as the nuclear norm (Candès et al.,, 2011; Nie et al.,, 2012; Klopp et al.,, 2017; Wong and Lee,, 2017; Chen et al.,, 2021). Under suitable assumptions on the matrices $L^{*}$, $S^{*}$ and on the set $\Omega$, these methods enjoy strong theoretical guarantees. However, these schemes are in general computationally slow, and do not scale well to large matrices. A different approach is to strictly enforce a rank $r$ solution. This can be done by optimization over the manifold of rank $r$ matrices (Yan et al.,, 2013; Cambier and Absil,, 2016), or by projecting matrices onto it. Other works factorize the target matrix as $L=UV^{\top}$, and optimize over the two factor matrices $U\in\mathbb{R}^{n_{1}\times r}$ and $V\in\mathbb{R}^{n_{2}\times r}$ (Yi et al.,, 2016; Lin et al.,, 2017; Zeng and So,, 2017; Huang et al.,, 2021; Wang et al.,, 2023). The resulting problem involves only $(n_{1}+n_{2})r$ variables instead of the $n_{1}\times n_{2}$ entries in $L$. Factorization based methods are thus in general more scalable and able to handle larger matrices.
Regarding the second issue, some works attenuate the effect of the outliers either by estimating the locations of the corrupted entries and removing them, or by estimating their values $S_{ij}$ for $(i,j)\in\Omega$ (Yan et al.,, 2013; Cherapanamjeri et al.,, 2017; Huang et al.,, 2021; Chen et al.,, 2021; Wang and Wei,, 2024). Other works mitigate the influence of outliers by using a robust norm on the difference $X-L$ in the objective function (Nie et al.,, 2012; Cambier and Absil,, 2016; Lin et al.,, 2017; Zhao et al.,, 2016; Zeng and So,, 2017; Wang et al.,, 2023).
@@
[…]
These shortcomings raise the following challenge: Develop a computationally efficient robust matrix completion method that can handle ill-conditioned matrices, overparameterization and requires only a small number of observed entries.
@@
In this work we make a step towards resolving this challenge. We propose a novel RMC method, denoted RGNMR, which overcomes the above limitations. Our proposed method is a simple iterative algorithm based on Gauss–Newton linearization and removal of entries suspected to be outliers. As seen in Figures 1, 2 and 3, empirically the performance of RGNMR is not affected by an overparameterized rank or by the condition number. Furthermore, RGNMR recovers underlying matrices with significantly fewer observed entries than existing methods. In Appendix B we show that RGNMR continues to outperform other methods under broader conditions, including both outliers and additive noise, non-uniform sampling and various other ranks, fraction of outliers, etc. Furthermore, we show that RGNMR perform well on a real data from computer vision, involving background extraction. In addition to its excellent empirical performance, in Section 3 we derive theoretical recovery guarantees for RGNMR. Specifically, we present two theorems. Theorem 3.1 states that given a suitable initialization, RGNMR recovers the target matrix at a linear rate. This theorem holds under the weakest known assumptions for factorization based RMC methods. Theorem 3.2 establishes that the initialization scheme we propose yields a suitable initial estimate.
@@ 2 The RGNMR algorithm
  • $\{X_{i,j}\mid(i,j)\in\Omega\}$ observed entries
  • $r$ rank of $L^{*}$
  • $k$ upper bound on the number of corrupted entries
  • $\begin{pmatrix}U_{0}\\ V_{0}\end{pmatrix}\in\mathbb{R}^{n_{1}+n_{2}\times r}$ initialization
  • $\Lambda_{0}$ initial estimate of the set of corrupted entries
  • $T$ maximal number of iterations
@@
$\mathcal{P}_{\Omega}(A)_{i,j}=\begin{cases}A_{i,j}&(i,j)\in\Omega,\\ 0&(i,j)\not\in\Omega.\\ \end{cases}$
@@
The optimization variables of RGNMR are two factor matrices $U\in\mathbb{R}^{n_{1}\times r},V\in\mathbb{R}^{n_{2}\times r}$ and a subset $\Lambda\subset\Omega$ which estimates the locations of the corrupted entries. RGNMR receives as input the following quantities; the subset of observed entries $\{X_{i,j}\mid(i,j)\in\Omega\}$; the rank of the target matrix $r$; an upper bound $k$ on the number of corrupted entries; initial values of factor matrices $(U_{0},V_{0})$, and of the set of corrupted entries $\Lambda_{0}\subset\Omega$.
At each iteration RGNMR performs the following two steps: (i) given the current set of suspected outlier entries $\Lambda$, update $U,V$ using the remaining entries $\mathcal{P}_{\Omega\setminus\Lambda}(X)$; (ii) given the updated matrix $L$, recompute the new set of suspected outliers $\Lambda$, by the $k$ entries with largest magnitude in $\mathcal{P}_{\Omega}(X-L)$.
Let us provide some motivation for the above two steps. If the estimated set of non corrupted entries, $\Omega\setminus\Lambda$, is large enough and contains only non-corrupted entries, namely $\Lambda_{*}\subset\Lambda$, then employing any vanilla matrix completion method would recover $L^{*}$ exactly from $\mathcal{P}_{\Omega\setminus\Lambda}(X)=\mathcal{P}_{\Omega\setminus\Lambda}(L^{*})$. Regarding the second step, consider an ideal case where at some iteration after step (i) RGNMR obtained $L=L^{*}$. Then at all the non-corrupted entries $(i,j)\in\Omega\setminus\Lambda_{*}$, the residual $X_{i,j}-L_{i,j}=0$ and $L^{*}$ is a fixed point of the algorithm. Therefore, given $L$ it is reasonable to estimate the corrupted entries by the entries with largest magnitude in $\mathcal{P}_{\Omega}(X-L)$. Iterating this process we hope to identify a sufficiently large subset of $\Omega\setminus\Lambda_{*}$ and consequently recover $L^{*}$.
We now describe the two steps of RGNMR. The first step builds upon the matrix completion approach of Zilber and Nadler, (2022). Formally, given the estimate $\Lambda_{t}$ of $\Lambda_{*}$ and the current factorization $(U_{t},V_{t})$, the updated matrices $\left(U_{t+1}V_{t+1}\right)$ are the solution of the following least squares problem,
@@
This problem is rank deficient with an infinite number of solutions. As in (Zilber and Nadler,, 2022), we choose the solution with minimal norm of the new estimate $\|U_{t+1}\|_{F}^{2}+\|V_{t+1}\|_{F}^{2}$.
@@
RGNMR iterates these two steps until convergence or until a maximal number of iterations $T$ is reached. Algorithm 1 presents an outline of RGNMR.
@@
The GNMR method of Zilber and Nadler, (2022) iterates Eq. (1) till convergence. If $\Omega\setminus\Lambda$ is large enough and contains only non corrupted entries, then after a few iterations GNMR would exactly recover $L^{*}$. However, since in the first few iterations the set $\Omega\setminus\Lambda$ often still includes a few corrupted entries, in our approach we re-estimate $\Lambda$ after each update of $(U_{t},V_{t})$ in (1). This yields considerable savings in terms of runtime.
@@ 2.1 Estimating the number of corrupted entries
One of the input parameters to RGNMR is the upper bound $k$ on the number of corrupted entries. At each iteration RGNMR removes the $k$ entries with largest residual error, and performs low-rank matrix completion on the remaining entries. For RGNMR to succeed, it is crucial for $k$ to be a tight upper bound. Indeed, if $k\gg k^{*}$ then too many entries are removed and on the remaining entries the corresponding matrix completion problem is ill posed.
@@
As the true number of corrupted entries $k^{*}$ is unknown, it is important to develop a method to tightly upper bound it. Here we propose a scheme to do so. Our approach is motivated by the empirical observation that the estimates $\Lambda_{t}$ of RGNMR behave differently if $k\leq k^{*}$ or $k>k^{*}$. If $k\leq k^{*}$ then the estimates $\Lambda_{t}$ of the set of corrupted entries converge after several iterations. However, if $k>k^{*}$, the sets $\Lambda_{t}$ do not converge. This is illustrated in Figure 4 in the appendix, and we provide some intuition for this behavior below. Building upon this observation, we propose a binary search algorithm to estimate $k^{*}$, assuming it belongs to the range $[k_{\min},k_{\max}]$. In practice, we may take $k_{\min}=0$ and $k_{\max}=|\Omega|/2$ (namely at most 50% corrupted entries). At each step we run RGNMR with $k=\lfloor(k_{\min}+k_{\max})/2\rfloor$. If the sets $\Lambda_{t}$ converged, then we update $k_{\min}=k$. Otherwise, we set $k_{\max}=k$. After $\log_{2}(k_{\max}-k_{\min})$ steps we obtain a tight upper bound on $k^{*}$.
@@
To provide intuition to why RGNMR behaves differently with $k\leq k^{*}$ or $k>k^{*}$, consider the following scenario. Assume that $k\leq k^{*}$ and that at some iteration, the set $\Lambda_{t}$ of $k$ entries with largest current residuals, satisfies that $\Lambda_{t}\subseteq\Lambda_{*}$. The next estimate $L_{t+1}$ of $L^{*}$ minimizes $\|\mathcal{P}_{\Omega\setminus\Lambda_{t}}(L-X)\|_{F}$, ignoring the entries in $\Lambda_{t}$, which by assumption are all outliers. Hence, $L_{t+1}$ tries to fit a low rank matrix to the remaining entries, and thus we expect the entries in $\Lambda_{t}$ to remain the $k$ largest residual entries. Therefore, $\Lambda_{t}=\Lambda_{t+1}$ and as the same argument holds for all subsequent iterations the sets $\Lambda_{t}$ converge. In contrast, if $k>k^{*}$ then at any iteration $t$, the set $\Lambda_{t}$ contains at least $k-k^{*}$ non-corrupted entries. Assume that at some iteration $t$, the method successfully detected all outliers, $\Lambda_{*}\subset\Lambda_{t}$. Then, if $k$ is not too large, the next estimate should be very close to the true low rank matrix $L_{t+1}\approx L^{*}$. As the computations are done in finite precision, the residuals at both $\Omega\setminus\Lambda_{t}$ as well as at $\Lambda_{t}\setminus\Lambda_{*}$ are not precisely zero but rather, due to rounding errors, appear as very small random values. Therefore, even though $\Lambda_{*}\subset\Lambda_{t+1}$ its remaining $k-k^{*}$ entries are not the same as those of $\Lambda_{t}\setminus\Lambda_{*}$, but rather change at each iteration $t$. Hence, the sets $\Lambda_{t}$ do not converge.
@@
Several RMC methods aim to remove the corrupted entries and require an estimate of $k^{*}$ (Yan et al.,, 2013; Yi et al.,, 2016). Since the intuition provided above holds also for these methods, a similar scheme to estimate $k^{*}$ can be incorporated there as well.

### v2

@@ Abstract
Recovering a low rank matrix from a subset of its entries, some of which may be corrupted, is known as the robust matrix completion (RMC) problem. Existing RMC methods have several limitations: they require a relatively large number of observed entries; they may fail under overparametrization, when their assumed rank is higher than the correct one; and many of them fail to recover even mildly ill-conditioned matrices. In this paper we propose a novel RMC method, denoted RGNMR, which overcomes these limitations. RGNMR is a simple factorization-based iterative algorithm, which combines a Gauss–Newton linearization with removal of entries suspected to be outliers. On the theoretical front, we prove that under suitable assumptions, RGNMR is guaranteed exact recovery of the underlying low rank matrix. Our theoretical results improve upon the best currently known for factorization-based methods. On the empirical front, we show via several simulations the advantages of RGNMR over existing RMC methods, and in particular its ability to handle a small number of observed entries, overparameterization of the rank and ill-conditioned matrices. In addition, we propose a novel scheme for estimating the number of corrupted entries. This scheme may be used by other RMC methods that require as input the number of corrupted entries.
@@ 1 Introduction
Figure 1: Performance of RMC methods as a function of the number of observed entries measured by the oversampling ratio $\frac{|\Omega|}{r\cdot(n_{1}+n_{2}-r)}$ where $r\cdot(n_{1}+n_{2}-r)$ is the number of degrees of freedom of a rank $r$ matrix. (left) Median rel-RMSE; (right) Failure probability ($\pm 1.96$ SE). The underlying matrix $L^{*}$ has a condition number $\kappa=2$. The corruption fraction is $\alpha=5\%$.
Figure 2: Performance of RMC methods under overparameterization with input rank of $5+i$ for $i\in[0,5]$. (left) Median rel-RMSE; (right) Failure probability ($\pm 1.96$ SE). The matrix $L^{*}$ has a condition number $\kappa=2$ and the oversampling ratio is $\frac{|\Omega|}{r\cdot(n_{1}+n_{2}-r)}=12$.The corruption fraction is $\alpha=5\%$
Figure 3: Performance of RMC methods as a function of the condition number. (left) Median rel-RMSE; (right) Failure probability ($\pm 1.96$ SE). The oversampling ratio is $\frac{|\Omega|}{r\cdot(n_{1}+n_{2}-r)}=12$. The corruption fraction is $\alpha=5\%$.
Low-rank matrices play a fundamental role in multiple scientific disciplines. As reviewed in Davenport and Romberg, (2016), in various applications, there is a need to recover a low rank matrix from only a subset of its entries. Examples include recommendation systems (Bennett et al.,, 2007), various problems in computer vision (Kennedy et al.,, 2016; Tulyakov et al.,, 2016; Miao and Kou,, 2021), in sensor networks (Wu et al.,, 2025) and in single‐cell data analysis (Lejun et al.,, 2025). A key challenge in these and other applications is that some of the observed entries may be arbitrarily corrupted outliers.
@@
Over the past decade, the RMC problem was studied from both mathematical, computational and statistical perspectives. On the computational front, several RMC algorithms have been proposed. In general, to solve the RMC problem the following optimization issues need to be addressed: (i) either promote a low rank solution or strictly enforce it; and (ii) suppress the potentially detrimental effect of the outliers, whose locations are a-priori unknown.
Regarding the first issue, one approach is to incorporate into the optimization objective a low-rank promoting penalty such as the nuclear norm (Candès et al.,, 2011; Nie et al.,, 2012; Klopp et al.,, 2017; Wong and Lee,, 2017; Chen et al.,, 2021). Under suitable assumptions on the matrices $L^{*}$, $S^{*}$ and on the set $\Omega$, these methods enjoy strong theoretical guarantees. However, these schemes are in general computationally slow, and do not scale well to large matrices. A different approach is to strictly enforce a rank $r$ solution. This can be done by optimization over the manifold of rank $r$ matrices (Yan et al.,, 2013; Cambier and Absil,, 2016), or by projecting matrices onto it. Several works factorize the target matrix as $L=UV^{\top}$, and optimize over the two factor matrices $U\in\mathbb{R}^{n_{1}\times r}$ and $V\in\mathbb{R}^{n_{2}\times r}$ (Yi et al.,, 2016; Lin et al.,, 2017; Zeng and So,, 2017; Huang et al.,, 2021; Wang et al.,, 2023). The resulting problem involves only $(n_{1}+n_{2})r$ variables instead of the $n_{1}\times n_{2}$ entries in $L$. Factorization based methods are thus in general more scalable and able to handle larger matrices.
Regarding the second issue, some works attenuate the effect of the outliers either by estimating the locations of the corrupted entries and removing them, or by estimating their values $S_{ij}$ for $(i,j)\in\Omega$ (Yan et al.,, 2013; Cherapanamjeri et al.,, 2017; Huang et al.,, 2021; Chen et al.,, 2021; Wang and Wei,, 2024). Other works mitigate the influence of outliers by using a robust norm on the difference XLX-L in the objective function (Nie et al.,, 2012; Cambier and Absil,, 2016; Lin et al.,, 2017; Zhao et al.,, 2016; Zeng and So,, 2017; Wang et al.,, 2023).
@@
We illustrate these issues in Figures 1, 2 and 3. These figures show recovery results of various RMC algorithms for a matrix $L^{*}$ of size $3200\times 400$, rank $r=5$, at a corruption rate of $\alpha=5\%$ (see Section 4 for further details). As in Zeng and So, (2017); Tong et al., (2021) and Huang et al., (2021), the quality of a recovered matrix $\hat{L}$ is measured by the relative error $\texttt{rel-RMSE}=\tfrac{\|\hat{L}-L^{*}\|_{F}}{\|L^{*}\|_{F}}$, and is considered a failure if $\texttt{rel-RMSE}>10^{-3}$. For each simulation setting we ran 100 independent realizations. As shown in these figures, RMC methods such as AOP (Yan et al.,, 2013), RMC (Cambier and Absil,, 2016), RPCA-GD (Yi et al.,, 2016), RRMC (Cherapanamjeri et al.,, 2017), HUB (Ruppel et al.,, 2020) and HOAT (Wang et al.,, 2023) require a large number of observed entries and fail when overparameterized. In addition, all of them (except HUB) fail at condition numbers greater than 5.
On the theoretical front, some RMC methods have no theoretical recovery guarantees. The currently available guarantees for other methods are somewhat limited. As detailed in Table 1, the guarantees for some methods only hold for fully observed matrices, or their allowed corruption level strongly depends on the matrix rank and condition number.
These shortcomings raise the following challenges: (i) Develop a computationally efficient RMC method able to handle ill-conditioned matrices, overparameterization, and a small number of observed entries; (ii) derive for this method strong recovery guarantees that improve upon those of existing methods.
@@
In this work we make a step towards resolving this challenge. We propose a novel RMC method, denoted RGNMR, which overcomes the above limitations. Our proposed method is a simple iterative algorithm based on Gauss–Newton linearization and removal of entries suspected to be outliers. As seen in Figures 1, 2 and 3, empirically RGNMR can successfully recover the underlying matrices with significantly fewer observed entries than existing methods. Furthermore, the performance of RGNMR is not affected by an overparameterized rank or by the condition number. In Appendix B we show that RGNMR continues to outperform other methods under broader conditions, including a combination of outliers and additive noise, non-uniform sampling, higher rank matrices, a range of fractions of outliers, etc. Furthermore, we show that RGNMR performs well on a real data set from computer vision, involving background extraction. In addition to its excellent empirical performance, in Section 3 we derive theoretical recovery guarantees for RGNMR. Specifically, we present two theorems. Theorem 3.1 states that given a suitable initialization, RGNMR recovers the target matrix at a linear rate. This theorem holds under the weakest known assumptions for factorization based RMC methods. Theorem 3.2 establishes that the initialization scheme we propose yields a sufficiently accurate initial estimate for Theorem 3.1 to hold.
@@ 2 The RGNMR algorithm
  • $\{X_{i,j}\mid(i,j)\in\Omega\}$ - observed entries
  • $r$ - rank of $L^{*}$
  • $k$ - upper bound on the number of corrupted entries
  • $\begin{pmatrix}U_{0}\\ V_{0}\end{pmatrix}\in\mathbb{R}^{(n_{1}+n_{2})\times r}$ - initialization
  • $\Lambda_{0}$ - initial estimate of the set of corrupted entries
  • $T$ - maximal number of iterations
@@
[…]
@@
The optimization variables of RGNMR are two factor matrices $U\in\mathbb{R}^{n_{1}\times r},V\in\mathbb{R}^{n_{2}\times r}$ and a subset $\Lambda\subset\Omega$ which estimates the locations of the corrupted entries. RGNMR receives as input the following quantities; the subset of observed entries $\{X_{i,j}\mid(i,j)\in\Omega\}$; the rank of the target matrix $r$; an upper bound $k$ on the number of corrupted entries; initial guess $(U_{0},V_{0})$ for the factor matrices and $\Lambda_{0}\subset\Omega$ for the set of corrupted entries.
At each iteration RGNMR performs the following two stpdf: (i) given the current set of suspected outlier entries $\Lambda$, update $U,V$ using the remaining entries $\mathcal{P}_{\Omega\setminus\Lambda}(X)$; (ii) given the updated matrix $L$, recompute the new set of suspected outliers $\Lambda$, by the $k$ entries with largest magnitude in $\mathcal{P}_{\Omega}(X-L)$.
Let us provide some motivation for the above two stpdf. If the estimated set of non corrupted entries, $\Omega\setminus\Lambda$, is large enough and contains only non-corrupted entries, namely $\Lambda_{*}\subset\Lambda$, then the matrix $L^{*}$ could be recovered exactly from $\mathcal{P}_{\Omega\setminus\Lambda}(X)=\mathcal{P}_{\Omega\setminus\Lambda}(L^{*})$ by solving a vanilla matrix completion problem, with no outliers or noise. Regarding the second step, consider an ideal case where after step (i) of some iteration RGNMR obtained $L=L^{*}$. Then at all the non-corrupted entries $(i,j)\in\Omega\setminus\Lambda_{*}$, the residual $X_{i,j}-L_{i,j}=0$ and $L^{*}$ is a fixed point of the algorithm. Therefore, given $L$ it is reasonable to estimate the corrupted entries by the entries with largest magnitude in $\mathcal{P}_{\Omega}(X-L)$. Iterating this process we hope to identify a sufficiently large subset of $\Omega\setminus\Lambda_{*}$ and consequently recover $L^{*}$.
We now describe the two stpdf of RGNMR. The first step builds upon the matrix completion approach of Zilber and Nadler, (2022). Formally, given the estimate $\Lambda_{t}$ of $\Lambda_{*}$ and the current factorization $(U_{t},V_{t})$, the updated matrices $\left(U_{t+1}V_{t+1}\right)$ are the solution of the following least squares problem,
@@
This problem is rank deficient with an infinite number of solutions. As in Zilber and Nadler, (2022), we choose the new estimate to be the solution with minimal norm $\|U_{t+1}\|_{F}^{2}+\|V_{t+1}\|_{F}^{2}$.
@@
RGNMR iterates these two stpdf until convergence or until a maximal number of iterations $T$ is reached. Algorithm 1 presents an outline of RGNMR.
@@
In general, removing some of the observed entries, as in (2), may be detrimental. For example, if after removal of entries suspected as outliers, only less than $r$ entries remain at a specific row or column, then the resulting matrix completion problem is ill-posed and exact recovery of $L^{*}$ is not possible. As illustrated empirically in Zilber and Nadler, (2022), GNMR is able to complete low rank matrices even if the number of observed entries is near the information limit. Therefore, as long as the number of remaining observed entries in each row and column is above $r$, our method is often still able to recover the target matrix.
Remark 2.2.
As shown in Zilber and Nadler, (2022), in the absence of outliers, GNMR which iteratively solves (1), can successfully recover ill-conditioned low rank matrices, from relatively few observed entries. However, as illustrated in Figure 5 in the appendix, GNMR is not robust and fails completely in the presence of even a small fraction of corrupted entries. Since in the first few iterations the set $\Omega\setminus\Lambda_{t}$ typically includes outliers, employing GNMR would not result in a significantly better update of $(U_{t},V_{t})$ than that of a single optimization step. Therefore, in our scheme we do not run GNMR till convergence. Instead we re-estimate $\Lambda$ after each update of $(U_{t},V_{t})$. This approach avoids redundant iterations of GNMR and yields considerable savings in terms of runtime.
@@ 2.1 Estimating the number of corrupted entries
One of the input parameters to RGNMR is $k$. At each iteration RGNMR removes the $k$ entries with largest residual error, and performs low-rank matrix completion on the remaining entries. For RGNMR to succeed, it is crucial for $k$ to be a tight upper bound of the true number of corrupted entries $k^{*}$. Indeed, if $k\gg k^{*}$ then too many entries are removed and on the remaining entries the corresponding matrix completion problem is ill posed. As the true number of corrupted entries $k^{*}$ is unknown, it is important to develop a method to tightly upper bound it. Here we propose a scheme to do so.
@@
Our approach is motivated by the empirical observation that the estimates $\Lambda_{t}$ of RGNMR behave differently if $k\leq k^{*}$ or $k>k^{*}$. If $k\leq k^{*}$ then the estimates $\Lambda_{t}$ of the set of corrupted entries converge after several iterations. However, if $k>k^{*}$, the sets $\Lambda_{t}$ do not converge. This is illustrated in Figure 4 in the appendix, and we provide some intuition for this behavior below. Building upon this observation, we propose a binary search algorithm to estimate $k^{*}$, assuming it belongs to the range $k_{\min},k_{\max}]$. In practice, we may take $k_{\min}=0$ and $k_{\max}=|\Omega|/2$ (namely at most 50% corrupted entries). At each step we run RGNMR with $k=\lfloor(k_{\min}+k_{\max})/2\rfloor$. If the sets $\Lambda_{t}$ converged, then we update $k_{\min}=k$. Otherwise, we set $k_{\max}=k$. After $\log_{2}(k_{\max}-k_{\min})$ stpdf we obtain a tight upper bound on $k^{*}$.
@@
To provide intuition to why RGNMR behaves differently with $k\leq k^{*}$ or $k>k^{*}$, consider the following scenario. Assume that $k\leq k^{*}$ and that at some iteration, the set $\Lambda_{t}$ of $k$ entries with largest current residuals, satisfies that $\Lambda_{t}\subseteq\Lambda_{*}$. The next estimate $L_{t+1}$ of $L^{*}$ minimizes $\|\mathcal{P}_{\Omega\setminus\Lambda_{t}}(L-X)\|_{F}$, ignoring the entries in $\Lambda_{t}$, which by assumption are all outliers. Hence, $L_{t+1}$ tries to fit a low rank matrix to the remaining entries, and thus we expect the entries in $\Lambda_{t}$ to remain the $k$ largest residual entries. Therefore, $\Lambda_{t}=\Lambda_{t+1}$ and as the same argument holds for all subsequent iterations the sets $\Lambda_{t}$ converge. In contrast, if $k>k^{*}$ then at any iteration $t$, the set $\Lambda_{t}$ contains at least $k-k^{*}$ non-corrupted entries. Assume that at some iteration $t$, the method successfully detected all outliers, i.e. $\Lambda_{*}\subset\Lambda_{t}$. Then, if $k$ is not too large, the next estimate should be very close to the true low rank matrix $L_{t+1}\approx L^{*}$. As the computations are done in finite precision, the residuals at both $\Omega\setminus\Lambda_{t}$ as well as at $\Lambda_{t}\setminus\Lambda_{*}$ are not precisely zero but rather, due to rounding errors, appear as very small random values. Therefore, even though $\Lambda_{*}\subset\Lambda_{t+1}$ its remaining $k-k^{*}$ entries are not the same as those of $\Lambda_{t}\setminus\Lambda_{*}$, but rather change at each iteration $t$. Hence, the sets $\Lambda_{t}$ do not converge.
@@
Several RMC methods aim to remove the corrupted entries and require an estimate of $k^{*}$ (Yan et al.,, 2013; Yi et al.,, 2016). Since the intuition provided above holds also for these methods, a similar scheme to estimate $k^{*}$ can be incorporated there as well. We note that in the simulations we conducted these methods were provided with the exact number of outliers. Still, in various settings, these methods failed to recover the underlying matrix, whereas RGNMR succeeded. This highlights that our novel scheme for estimating the number of outliers is not the sole advantage of our proposed method.